OpenStudy (abbysmith12):

I will medal and fan. Whats the product? 2/3 x 5/8 x 3/4 A. 5/16 B. 5/32 C. 7/12 D. 2/3

OpenStudy (freckles):

\[\frac{2}{3} \cdot \frac{5}{8} \cdot \frac{3}{4} \\ \text{ move numerators around } \frac{3}{3} \cdot \frac{2}{8} \cdot \frac{5}{4}\] what do you notice about 3/3?

OpenStudy (abbysmith12):

improper?

OpenStudy (freckles):

3 divided by 3 equals?

OpenStudy (abbysmith12):

@quin100 @Write0The0Right0For0Wrong0Doing @Ellia201 @Rea201 @TTP @yahman @umartouseef @iamabarbiegirl @ogjb_ @peterg2001 @adrynicoleb @sleepyhead314 @Destinymasha @flasht @Glori96 @Hotchellerae21 @justuu @KyanTheDoodle @lizzyrockefeller @xo.minnie.xox @Compassionate @VJT @Brostep0s @nevermind_justschool @marylou004

OpenStudy (justuu):

Wait a second